OpenStudy (anonymous):

Choose the replacement for x and y that makes this equation a true statement: 2x + 55y = −214 a. x = −3 and y = −4 b. x = −4 and y = 3 c. x = 3 and y = −4 d. x = 4 and y = 4

OpenStudy (anonymous):

wow that completly messed up :/

OpenStudy (anonymous):

2x + 55y = −214 2(0) + 55y = −214 55y = −214 -214 / 55 = y 2x + 55y = −214 2x + 55(0) = −214 2x = −214 -214 / 2 = x